# Brain-inspired sparse training enables Transformers and LLMs to perform as fully connected

### v1

@@ 4.4 Natural Language Generation
We utilize LLaMA-130M (Touvron et al., 2023a) architecture as the baseline for our language generation experiments. For LLaMA, we follow the experiment setup from (Zhao et al., 2024) and detailed in Table 7. To ensure the FLOPs are the same for all the density decrease methods, for CHTss, we implement a half-step strategy for the density decay steps.
As shown in Table 3, we show the validation perplexity results of each algorithm across the density decrease. As shown, CHTs stably outperformance SET and RigL while CHTss constantly better than GraNet and GMP. At 70% sparsity, CHTss is already able to perform a comparable performance in comparison to the fully connected. It is important to note that RigL and GraNet exhibit subpar performance in this evaluation due to the use of bfloat16 precision in this configuration. This lower precision adversely impacts gradient accuracy, particularly in the early stages of training. Since both RigL and GraNet rely on gradient information to regrow new links, the imprecise gradients lead to regrowing wrong links, thereby hindering their performance.
Zero-shot evaluations.
The pretrained model of CHTss with 30% sparsity and the fully connected one are evaluated the zero-shot performance on the GLUE (Wang et al., 2019)(cola, sst2, mrpc, qqp, mnli, qnli, rte, wnli) and SuperGLUE (Wang et al., 2020) (boolq, hellaswag, CB, copa). We show the results in Table 4. The performance difference between FC and CHTss is statistically significant (p-value = 0.01) according to a paired two-sided Wilcoxon signed rank test, which means CHTss is significantly better than fully connected ones in zero-shot evaluations.
Table 4: Zero-shot evaluation of LLaMA-130M between fully connected pertrained and pretrained with CHTss (70% sparsity) across GLUE and Superglue datasets. MRPC and QQP use F1 scores while the others use ACC. $s_{i}$ indicates the starting sparsity of the dst methods that use density decay strategy.
Dataset FC CHTss ($s_{i}=0.5$)
CoLA 65.29 ± 1.47 69.13 ± 1.43
MNLI 32.44 ± 0.47 32.72 ± 0.47
MRPC 64.96 ± 2.36 81.05 ± 1.64
QNLI 50.38 ± 0.68 49.37 ± 0.68
QQP 52.09 ± 0.28 53.82 ± 0.26
RTE 48.38 ± 3.01 50.54 ± 3.01
SST-2 49.54 ± 1.69 49.08 ± 1.69
WNLI 49.30 ± 5.98 52.11 ± 5.97
Hellaswag 26.95 ± 0.44 26.97 ± 0.44
Boolq 43.85 ± 0.87 56.79 ± 0.87
CB 46.43 ± 6.72 50.00 ± 6.74
Copa 56.00 ± 4.99 57.00 ± 4.98
AVG 48.80 52.38
Win rate 0.17 0.83
5 Conclusion and Discussion
We move forward the current knowledge in brain-inspired dynamic sparse training, proposing the Cannistraci-Hebb Training soft rule with sigmoid gradual density decay (CHTss). First, we introduce a matrix multiplication mathematical formula for GPU-friendly approximation of the CH link predictor. This significantly reduces the computational complexity of CHT and speeds up the running time, enabling the implementation of CHT in large-scale models. Second, we propose a Cannistraci-Hebb training soft rule (CHTs), which innovatively utilizes a soft sampling rule for both removal and regrowth links striking a balance for epitopological exploration and exploitation. Third, in transformer-based models, we integrate CHTs with a sigmoid gradual density decay strategy (CHTss). Empirically, CHTs demonstrate a remarkable ability to achieve ultra-sparse configurations—up to 99% sparsity in MLPs for image classification—surpassing fully connected networks. Notably, the regrowth process under CHTs does not rely on gradients. In contrast to RigL, which depends on inputs, topology, weights, and activation functions for predicting new connections, CHTs’ regrowth is merely brain-inspired and network topology-driven. With the sigmoid gradual density decay, CHTss surpasses the fully connected Transformer using only 5% density and achieves comparable language modeling performance, along with better zero-shot results, to the fully connected LLaMA-130M at just 30% density. This represents a relevant result for dynamic sparse training. We describe the limitations of this study and future works in Appendix A.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Limitation
A potential limitation of this work is that the hardware required to accelerate sparse training with unstructured sparsity has not yet become widely adopted. Consequently, this article does not present a direct comparison of training speeds with those of fully connected networks. However, several leading companies (Thangarasa et al., 2023) have already released devices that support unstructured sparsity in training.
For future work, we aim to develop methods for automatically determining the temperature for soft sampling at each epoch, guided by the topological features of each layer. This could enable each layer to learn its specific topological rules autonomously. Additionally, we plan to test CHTss in larger LLMs such as LLaMA-1b and LLaMA-7b to evaluate the performance in scenarios with denser layers.
Appendix B Broader Impact
Appendix C Sparse topological initialization
Correlated sparse topological initialization.
Correlated Sparse Topological Initialization (CSTI) is a physical-informed topological initialization. CSTI generates the adjacency matrix by computing the Pearson correlation between each input feature across the calibration dataset and then selects the predetermined number of links, calculated based on the desired sparsity level, as the existing connections. CSTI performs remarkably better when the layer can directly receive input information. However, for layers that cannot receive inputs directly, it cannot capture the correlations from the start since the model is initialized randomly, as in the case of the Transformer. Therefore, in this article, we aim to address this issue by investigating different network models to initialize the topology, with the goal of improving the performance for cases where CSTI cannot be directly applied.
Bipartite scale-free model.
In artificial neural networks (ANNs), fully connected networks are inherently bipartite. This article explores initializing bipartite networks using models from network science. The Bipartite Scale-Free (BSF) (Zhang et al., 2024b) network model extends the concept of scale-freeness to bipartite structures, making them suitable for dynamic sparse training. Initially, the BSF model generates a monopartite Barabási-Albert (BA) model (Barabási & Albert, 1999), a well-established method for creating scale-free networks in which the degree distribution follows a power law ($\gamma$=2.76 in Figure 3). Following the creation of the BA model, the BSF approach removes any connections between nodes of the same type (neuron in the same layer) and rewires these connections to nodes of the opposite type (neuron in the opposite layer). This rewiring is done while maintaining the degree of each node constant to preserve the power-law exponent $\gamma$.
Bipartite small-world model.
The Bipartite Small-World (BSW) network model (Zhang et al., 2024b) is designed to incorporate small-world properties and high clustering coefficient into bipartite networks. Initially, the model constructs a regular ring lattice and assigns two distinct types of nodes to it. Each node is connected by an equal number of links to the nearest nodes of the opposite type, fostering high clustering but lacking the small-world property. Similar to the Watts-Strogatz model (WS) (Watts & Strogatz, 1998), the BSW model introduces a rewiring parameter, $\beta$, which represents the percentage of links randomly removed and then rewired within the network. At $\beta=1$, the model transitions into an Erdős-Rényi model (ERDdS & R&wi, 1959), exhibiting small-world properties but without high clustering coefficient, which is popular as the topological initialization of the other dynamic sparse training methods (Mocanu et al., 2018; Evci et al., 2020; Yuan et al., 2021).
Figure 4: The ablation test of the $\beta$ of the bipartite small world model and the removal methods in CHTs. a) evaluates the influence of the rewiring rate $\beta$ on the model performance when initialized with the Bipartite Small-World network model (BSW). b) assess the influence of link removal selecting from the weight magnitude (WM), weight magnitude soft (WM-soft), relative importance (RI), and relative importance soft (RI-soft). We utilize the win rate of the compared factors under the same setting across each realization of 3 seeds for all experiment combinations on MLP. The factor with the highest win rate is highlighted in orange.
Appendix D Equal Partition and Neuron Resorting to enhance bipartite scale-free network initialization
As indicated in SET and CHT (Mocanu et al., 2018; Zhang et al., 2024b), trained sparse models typically converge to a scale-free network. This suggests that initiating the network with a scale-free structure might initially enhance performance. However, starting directly with a Bipartite Scale-Free model (BSF, power-law exponent $\gamma=2.76$) does not yield effective results. Upon deeper examination, two potential reasons emerge:
  • As demonstrated in CHT, in the final network, the hub nodes of one layer’s output should correspond to the input layer of the subsequent layer, which means the hub nodes should have a high degree on both sides of the layer. However, the BSF model’s random selection disrupts this correspondence, significantly reducing the number of Credit Assignment Paths (CAP) (Zhang et al., 2024b) in the model. CAP is defined as the chain of the transformation from input to output, which counts the number of links that go through the hub nodes in the middle layers.
To address these issues, we propose two solutions:
Figure 5: The Performance of the bipartite scale-free model and two enhanced techniques. a) shows the win rate of the Bipartite Scale-Free network model (BSF) with the different techinques. EP stands for equal partition of the first layer, and Resort refers to reordering the neurons based on their degree. b) assesses the comparison between Correlated Sparse Topological Initialization (CSTI), the Bipartite Scale-Free (BSF) model with the best solution from a), and the Bipartite Small-World (BSW) model with $\beta=0.25$.
Appendix E Network percolation and extension to Transformer.
We have adapted network percolation (Li et al., 2021; Zhang et al., 2024b) to suit the architecture of the Transformer after link removal. The underlying concept involves identifying inactive neurons, which we define as those lacking connections on one or both sides of a neuron layer. Such neurons disrupt the flow of information during forward propagation or backpropagation. In addition, Layer-wise computation of the CH link prediction score further implies that neurons without connections on one side are unlikely to form connections in the future. Therefore, network percolation becomes essential to optimize the use of remaining links.
As shown in Figure 1, network percolation encompasses two primary processes: c1) inactive neuron removal to remove the neurons that lack connections on one or both sides; c2) incomplete path adjustment to remove the incomplete paths where links connect to the inactive neurons after c1). Typically applied in simpler continuous layers like those in an MLP, network percolation requires modification for more complex structures. For example, within the Transformer’s self-attention module, the outputs of the query and key layers undergo a dot product operation. It necessitates percolation in these layers to examine the activity of the neurons in both output layers at the same position. Similar interventions are necessary in the up_proj and gate_proj layers of the MLP module in the LLaMA model family (Touvron et al., 2023a, b).
Appendix F Baseline Methods
F.1 Fixed Density Dynamic Sparse Training Methods
SET
(Mocanu et al., 2018): Removes connections based on weight magnitude and randomly regrows new links.
RigL
(Evci et al., 2020): Removes connections based on weight magnitude and regrows links using gradient information, gradually reducing the proportion of updated connections over time.
CHT
(Zhang et al., 2024b): A state-of-the-art (SOTA) gradient-free method that removes links with weight magnitude and regrows links based on CH3-L3 scores. (Note: CHT is only evaluated on MLPs due to its computational cost in large models.)
F.2 Gradual Density Decrease Dynamic Sparse Training Methods
GMP
(Han et al., 2015; Zhu & Gupta, 2017): Prunes the network with weight magnitude and gradually decrease the density based on Equation 8. Although originally a pruning method, GMP is treated as a dynamic sparse training method in their implementation (Zhu & Gupta, 2017), as it stores historical weights and allows pruned weights to reappear during training since, during training, the pruning threshold might change.
MESTEM&S
(Yuan et al., 2021): Implements a two-stage density decrease strategy as described in the original work. It removes links based on the combination of weight magnitude and 0.01*gradient and regrows new links randomly.
GraNet
(Liu et al., 2021): Gradually decreases density using Equation 8. Similar to RigL, GraNet removes links based on the weight magnitude and regrows new links with the gradient of the existing links.

### v2

@@ Appendix I Integral of Sigmoid Density Decay and Cubic Density Decay
Table 14: Validation perplexity of different dynamic sparse training (DST) methods on OpenWebText using LLaMA-1B across varying sparsity levels.. Lower perplexity corresponds to better model performance. The performances that surpass the fully connected model are marked with “*".
Sparsity 0.7 0.9 0.95
FC 14.62
CHTs 14.53* 17.14 18.93
CHTss 15.15 15.62 16.51
We present a comparison of CHTs, CHTss, and fully connected network on language modeling tasks using the LLaMA-1B model on Table 14. The results clearly demonstrate that CHTs consistently outperform the fully connected (FC) baseline at 70%, even at a high sparsity of 95%, CHTss achieves a perplexity of 16.51, which is remarkably close to the FC baseline.
Zero-shot performance.
We also compare the performance of CHTs, CHTss, and the fully connected network on zero-shot tasks, as summarized in Table 15. The results demonstrate that CHTs generally outperforms the fully connected baseline at 90% and 95% sparsity levels. This suggests that CHTs enhances the model’s generalization capability on previously unseen data.
Table 15: Zero-shot evaluation of LLaMA-1B across GLUE and SuperGLUE. ACC scores are shown. Values are mean ± sd over 5 seeds (lm-eval). Red values indicate the top scores over models and sparsities for a benchmark.
Dataset FC 70 % sparsity 90 % sparsity 95 % sparsity
CHTs CHTss CHTs CHTss CHTs CHTss
CoLA 40.27 ± 1.52 38.83 ± 1.51 44.20 ± 1.54 50.53 ± 1.55 67.88 ± 1.45 68.84 ± 1.43 31.83 ± 1.44
MNLI 32.89 ± 0.47 32.77 ± 0.47 32.65 ± 0.47 32.83 ± 0.47 32.73 ± 0.47 32.91 ± 0.47 32.68 ± 0.47
MRPC 39.95 ± 2.43 40.44 ± 2.43 36.03 ± 2.38 66.42 ± 2.34 38.48 ± 2.41 67.65 ± 2.32 46.81 ± 2.47
QNLI 49.95 ± 0.68 50.96 ± 0.68 49.70 ± 0.68 50.36 ± 0.68 49.79 ± 0.68 49.26 ± 0.68 51.42 ± 0.68
QQP 47.94 ± 0.25 43.40 ± 0.25 49.12 ± 0.25 50.57 ± 0.25 48.22 ± 0.25 36.92 ± 0.24 41.70 ± 0.25
RTE 47.29 ± 3.01 46.93 ± 3.00 55.96 ± 2.99 46.57 ± 3.00 52.71 ± 3.01 52.35 ± 3.01 49.46 ± 3.01
SST-2 65.71 ± 1.61 52.98 ± 1.69 49.66 ± 1.69 49.08 ± 1.69 53.21 ± 1.69 54.59 ± 1.69 49.08 ± 1.69
WNLI 50.70 ± 5.98 49.30 ± 5.98 47.89 ± 5.97 56.34 ± 5.93 50.70 ± 5.98 50.70 ± 5.98 40.85 ± 5.88
Hellaswag 28.98 ± 0.45 28.75 ± 0.45 28.70 ± 0.45 27.57 ± 0.45 28.13 ± 0.45 27.57 ± 0.45 27.53 ± 0.45
Boolq 41.07 ± 0.86 50.89 ± 0.87 48.62 ± 0.87 44.59 ± 0.87 47.46 ± 0.87 55.38 ± 0.87 52.14 ± 0.87
CB 48.21 ± 6.74 50.00 ± 6.74 50.00 ± 6.74 50.00 ± 6.74 50.00 ± 6.74 37.50 ± 6.53 48.21 ± 6.74
Copa 65.00 ± 4.79 62.00 ± 4.88 64.00 ± 4.82 66.00 ± 4.76 63.00 ± 4.85 60.00 ± 4.92 60.00 ± 4.92
Average 46.50 45.60 46.38 49.24 48.53 49.47 44.31
Appendix M Experiments compute resources
All experiments were conducted on NVIDIA A100 80GB GPUs. MLP and Transformer models were trained using a single GPU, while LLaMA models were trained using eight GPUs in parallel.